OpenStudy (crisulcampo):

what is the implicit derivate dy/dx about this expression: (bx)^2 +(ay)^2 = (ab)^2 please folks!! help me

OpenStudy (crisulcampo):

@zepdrix

OpenStudy (crisulcampo):

@mathstudent55

OpenStudy (sshayer):

\[2 bx*b+2ay*a\frac{ dy }{ dx }=0\] \[2a^2y\frac{ dy }{ dx }=-2b^2x\] \[\frac{ dy }{ dx }=-\frac{ b^2x }{ a^2y }\]

OpenStudy (crisulcampo):

what about \[d^{2}y/dx^2\]

OpenStudy (crisulcampo):

@zepdrix @Zeronknight @ShadowLegendX

zepdrix (zepdrix):

Hold onto your first derivative, just keep it in your back pocket for now.\[\large\rm \color{indianred}{y'=\frac{-b^2x}{a^2y}}\]We'll need it later. From here, how bout applying your quotient rule, ya?\[\large\rm y''=\left(\frac{-b^2x}{a^2y}\right)'\]

zepdrix (zepdrix):

\[\large\rm y''=\frac{(-b^2x)'(a^2y)-(b^2x)(a^2y)'}{(a^2y)^2}\]